OpenStudy (anonymous):

Anyone here good with ellipses? If so, please help!! Find the center, vertices, and foci of the ellipse with equation x squared divided by four hundred plus y squared divided by two hundred and fifty six = 1 A. Center: (0, 0); Vertices: (-20, 0), (20, 0); Foci: (-12,0), (12, 0) B. Center: (0, 0); Vertices: (-20, 0), (20, 0); Foci: (-16,0), (16, 0) C. Center: (0, 0); Vertices: (0, -20), (0, 20); Foci: (0,-16), (0, 16) D. Center: (0, 0); Vertices: (0, -20), (0, 20); Foci: (0,-12), (0, 12) I know the answer is either B or C

OpenStudy (anonymous):

And also for Find an equation in standard form for the ellipse with the vertical major axis of length 6, and minor axis of length 4. I am pretty sure the answer is: x^2/9 + y^2/4 = 1 Is that correct?

Nnesha (nnesha):

I have few minutes lets see if i can help u :=)

Nnesha (nnesha):

1st) is it vertical or horizontal ?

OpenStudy (anonymous):

Ok thank you! And for the first one it is horizontal. @Nnesha

Nnesha (nnesha):

yes right so formula for foci \[\huge\rm (h \pm c , k)\]

Nnesha (nnesha):

first of all you need to find c

Nnesha (nnesha):

use this equation \[\huge\rm c^2=a^2-b^2\]

OpenStudy (anonymous):

Would c = SqRt656? @Nnesha

Nnesha (nnesha):

remember (h,k) is the center point it's horizontal so only x values would change y would stay same no change in y values

Nnesha (nnesha):

no what is a^2 and b^2 ?

Nnesha (nnesha):

\[\huge\rm \frac{ (x-h)^2 }{ a^2 }+\frac{ (y-k)^2 }{ b^2 }\] standard form of ellipse

OpenStudy (anonymous):

Wouldn't a^2 = 256 and b^2 = 400?

OpenStudy (anonymous):

Oh wait.

OpenStudy (anonymous):

12 = c?

Nnesha (nnesha):

okay for ellipse a= bigger number if a is under the x then it's horizontal and if a is under y then it would be vertical

Nnesha (nnesha):

yes !

Nnesha (nnesha):

so now what are the foci points ?

OpenStudy (anonymous):

+/12, 0?

OpenStudy (anonymous):

Forgot the ( ) sorry

Nnesha (nnesha):

\[(\pm 12,0)\] looks great should be plus minus sign

OpenStudy (anonymous):

Yes that's what I meant, sorry.

OpenStudy (anonymous):

So my answer would actually be A. Ok, thanks! Could you help me with the 2nd one?

OpenStudy (anonymous):

And also with this one: The rectangular piece of property measures 10 mi by 4 mi. Find an equation for the ellipse if the path is to touch the center of the property line on all 4 sides. I believe my answer would be: x^2/25 + y^2/4 = 1 But I just wanted to make sure.

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @jdoherty And also for Find an equation in standard form for the ellipse with the vertical major axis of length 6, and minor axis of length 4. I am pretty sure the answer is: x^2/9 + y^2/4 = 1 Is that correct? \(\color{blue}{\text{End of Quote}}\) *vertical*

OpenStudy (anonymous):

? @Nnesha

Nnesha (nnesha):

if ellipse is vertical then bigger number supposed to be under x or y ?

OpenStudy (anonymous):

Y?

Nnesha (nnesha):

yes right try to draw points on the paper

Nnesha (nnesha):

gtg sorry. :(

OpenStudy (anonymous):

Ok :( Are you free at all tomorrow?

OpenStudy (anonymous):

@saseal Could you help with the other two questions?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank you! @saseal

OpenStudy (anonymous):

1. The rectangular piece of property measures 10 mi by 4 mi. Find an equation for the ellipse if the path is to touch the center of the property line on all 4 sides. I believe my answer would be: x^2/25 + y^2/4 = 1 But I just wanted to make sure. 2. Find an equation in standard form for the ellipse with the vertical major axis of length 6, and minor axis of length 4. I am pretty sure the answer is: x^2/9 + y^2/4 = 1 Is that correct?

OpenStudy (anonymous):

@saseal ^

OpenStudy (anonymous):

first one is correct...2nd one is debatable because the question is not very specified

OpenStudy (anonymous):

if they specifed the major axis is x and minor is y then you are correct

OpenStudy (anonymous):

Hmm...how could I determine for sure?

OpenStudy (anonymous):

hmm you are wrong...it says vertical major axis

OpenStudy (anonymous):

Oh ok. So that would make the y axis the major axis?

OpenStudy (anonymous):

Oh wait. would it be x^2/4 + y^2/9 = 1? @saseal

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Could you help me a couple more questions? I kind of am just getting this stuff. Lol. @saseal

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation x squared over nine minus y squared over sixteen = 1 This is for Hyperbolas...

OpenStudy (anonymous):

\[\frac{ x^2 }{ 3^2 } - \frac{ y^2 }{ 4^2 } = 1\]

OpenStudy (anonymous):

now it becomes y^2/4^2 - x^2/3^2

OpenStudy (anonymous):

\[\frac{ y^2 }{ 4^2 } - \frac{ x^2 }{ 3^2 } =1\]

OpenStudy (anonymous):

Ok so then what can I do?

OpenStudy (anonymous):

you know where thid hyperbola opens open?

OpenStudy (anonymous):

Not really like when it's graphed?

OpenStudy (anonymous):